OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3. A. –52 B. –21 C. –22 D. –15

OpenStudy (anonymous):

Just add the two functions..

OpenStudy (anonymous):

yes i get that but could you write out what i have to do please im really sorry for the trouble

OpenStudy (anonymous):

Wait, then..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have two functions, f(x) and g(x).. Question is : you have to find f(x) + g(x) at x = -3

OpenStudy (anonymous):

Now, adding two functions is just similar to the way you add two numbers like 3 + 4 = 7

OpenStudy (anonymous):

So: \[f(x ) + g(x) = 4x - 7 + 2x + 4 \implies ??\]

OpenStudy (anonymous):

-12-7+-6+4? is what i have to do correct

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

so f(-3)+g(-3)=-21

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

You can add them first and then put the values also... :)

OpenStudy (anonymous):

-21 is correct..

OpenStudy (anonymous):

Or you can also do like : f(-3) + g(-3).. All are one and the same thing..

OpenStudy (anonymous):

ok Thank you very much

OpenStudy (anonymous):

tkhunny will you help me with another

OpenStudy (anonymous):

If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x)

OpenStudy (tkhunny):

Why is there still a question? f(-3) = 4(-3) - 7 = -12 - 7 = -19 g(-3) = 2(-3) + 4 = -6 + 4 = -2 f(-3) + g(-3) = -19 + (-2) = -19 - 2 = -21 OR h(x) = f(x) + g(x) = 4x – 7 + 2x + 4 = 6x - 3 h(-3) = 6(-3) - 3 = -18 - 3 = -21 If you've other problems, please post a new thread. ALWAYS show your own work.

OpenStudy (anonymous):

i hate functions